OpenStudy (anonymous):

the vertex of the porabla y=(x+8)^2-64 is located at A.(0,0) B.(-8,-64) C.(-8,0) D.(8,64)

OpenStudy (anonymous):

@mathmale can u help

OpenStudy (tkhunny):

Eyeball problem. You should recognize the "Vertex Form". (-8,-64) Done!

OpenStudy (anonymous):

can u help with more

OpenStudy (tkhunny):

Not if they are like this. That is ALL there is to know. Vertex Form? Read the vertex.

OpenStudy (anonymous):

can u help with ruadratic equations

OpenStudy (anonymous):

quadratic

OpenStudy (anonymous):

which of the following quadratic equations have the solutions x=-2 and x=5 A.-1/13(x-2)(x+5)=0 B.(x-2)(x+5)=0 C.(x+2)(x-5)=0 D.-1/13(x+2)(x-5)=0

OpenStudy (tkhunny):

Post on a new thread.